### v1

@@ 5 Trust Anchors and Assumptions for Control-Flow Attestation
The trust anchor is commonly used to protect $\mathcal{P}$’s measurement procedure and any key material used in generating attestation reports. For some proposals, this requires that the trust anchor passes traditional static attestation and that it has dedicated memory to store material that can only the anchor itself can access [41, 37, 13, 34]. For proposals that rely on TEEs, it is assumed that potentially applicable high-cost attacks cannot be performed by an adversary, such as hardware fault injections and side-channel attacks [50]. TEE-based proposals sometimes place trust in certain operating system components, such as kernel modules that initiate the communication with the TEE [38, 6]. In proposals where the kernel serves as $\mathcal{P}$’s trust anchor, it is somtimes assumed that a method for protecting cryptography keys is available, such as a TPM or a proprietary facility (see Intel MPK in [58]).
@@
[…]
@@ 5.3 Central Processing Unit Features
Another common trust anchor is the use of custom or proprietary or CPU features. These are predominately used to realise a measurement module that gathers control-flow information while the target program is being executed, i.e. at a machine code level. Such CFA schemes can be divided into two sub-categories. The first comprises schemes that employ custom extensions to CPU designs, e.g. open-source soft-core CPUs, to capture relevant control-flow events. The second contains proposals that rely on proprietary features of existing CPUs, such as tracing modules that track the history of instructions executed by the processor.
@@ 5.3.1 Custom Extensions
Figure 11: An overview of Atrium’s system architecture [37].
@@
Atrium [37] (Figure 11) uses hardware extensions to capture instructions and memory addresses at the execution stage of the CPU’s pipeline. The extensions capture the control-flow changes taken by the target program on $\mathcal{P}$ while executing a challenge from $\mathcal{V}$. An instruction filter is used to consider only, and all, branch and jump instructions in addition to the current program counter values, and a separate loop encoder is used to track the execution of loops. A hashing engine is used to convert the executed instruction and loop and function IDs—corresponding to nodes of a CFG—into a sequence of cryptographic hashes produced in real-time during the program’s execution. These hashes are continuously stored in memory and used in the attestation report returned to $\mathcal{V}$. It is noted that Atrium has reduced on-chip memory requirements while also addressing TOCTOU attacks when compared with LO-FAT.
LiteHAX [34] is more complicated in nature, hooking three parts of the CPU’s pipeline: the decode, execute, and load/store phases (Figure 13). The proposal monitors all control-flow instructions and memory addresses of load/store instructions in order to address control- and data-flow attacks respectively. A similar hash-based measurement approach is used as Atrium [37] and LO-FAT [12]. However, LiteHAX the hash engine takes encoded control-flow instructions and load/store instructions, their memory addresses, and the current program counter value when calculating the hash values for the attestation report.
Figure 12: System architecture of LO-FAT [12].
Figure 13: System architecture of LiteHAX [34].
@@ 5.3.2 Trace Modules
In recent years, CPU manufacturers have deployed capabilities that actively record control-flow information of programs under execution. This is known as program tracing whereby the CPU outputs encoded control-flow information into readable memory. Authorised applications can read this data, also known as a trace, to learn the control-flow path followed by a given program. The trace feature is supported by many contemporary CPU processors; for example, Intel CPUs support a tracing module from their 5th generation models, called Intel Processor Trace (PT) [64], which is designed for debugging pruposes. ARM CoreSight [65] is a comprehensive suite of debug and trace components that are embedded in ARM-based processors, providing developers with method to optimise and debug their applications on ARM architectures. The RISC-V specifications also describe a trace module; however, its availability varies according to vendor in the spirit of RISC-V’s modular instruction set architecture [66, 67].
@@ 5.4 Discrete Hardware Modules
Geden & Rasmussen [47] proposed an on-SoC hardware security module (HSM) that collects control-flow information from the system bus (Figure 14). The HSM continuously takes control-flow events as input from the CPU, including the requested instruction address, opcode, and operand values over the system-on-chip’s control line and address and data buses. The collected events are compared with security policies defined in the HSM’s runtime integrity model for detecting code injection and re-use attacks. The use of continuous run-time data addresses TOCTOU attacks: a challenge-response protocol that measures events at only fixed points in time does not detect attackers that act transiently between those points. The proposal allows $\mathcal{V}$ to acquire a report from the HSM showing whether any security policy violations were detected; it also supports local verification whereby the module outputs a binary result to indicate if a target control-flow hijacking attacks are detected. In scenarios where attacks are detected, the module outputs a binary result and provides auxiliary information, such as the last executed instruction.
Figure 14: Block diagram of the HSM used by Geden and Rasmussen (from [47]).
@@ 6 Instrumentation Methods
The collection of control-flow data can be achieved through the process of instrumentation, which involves modifying the target program to add instructions that capture, and measure, control-flow events of interest. Dispatch instructions are used near target control-flow transfers or at the entry of BBLs to this end. When a target control-flow transfer is executed at run-time, the dispatch instructions call pass the transfer(s) to the measurement module for it to be measured. Execution is resumed after the measurement process has completed.
@@ 6.2 Binary Analysis and Rewriting
Recording every function call can impose significant performance overhead (see p. 3, [58]). To address this, ReCFA [58] derives a CFG from the target program using TypeArmor for performing call-site filtering, whereby CFG nodes are skipped where none of its predecessors has more than one successor. ReCFA uses Dyninst [81] for the static binary instrumentation of user-space programs on $\mathcal{P}$. Dyninst provides a set of libraries and utilities designed for the static (and dynamic) inspection and manipulation of executables. Dyninst enables users to insert, modify, or remove code in a program’s CFG without access to its source code. Intel’s MPK system is used to protect the integrity of instrumented data structures—the loop stack and path stack—used by control-flow events. The authors compare this to SCaRR [13] where control-flow edges are measured in kernel space, which has a “potential enormous context switching cost for the control-flow events on an order of magnitude in gigabytes” (p. 6, [58]). The authors combine a kernel mode-based functions to preserve code integrity and MPK-protect data structures in order to uphold the integrity of the measurement procedure.
@@ 6.3 Dynamic Instrumentation
Dynamic instrumentation refers to the method of inserting monitoring and verifying code into a program while it is running. This compares to instrumenting target programs using static analysis and manipulating the executable before execution (e.g. using binary rewriting). Dynamic methods rely on the insertion of probes or tracking mechanisms into a running program. These probes can monitor the control flow of the application in real-time, identifying any deviations from expected paths.
@@ 7 Measuring and Monitoring Control Flow
The measurement process is one of the key facets of a CFA scheme. It is at the core of $\mathcal{P}$ demonstrating to $\mathcal{V}$ that an authorised path/walk was followed through the target program’s CFG. For a given CFA scheme, the measurement method typically remains constant, regardless of the control-flow paths or the target program in question. However, these methods tend to vary significant between CFA schemes, which we elaborate upon in this section.
@@
A CFA scheme may measure the whole program and partial elements, or multiple aspects of partial elements, simultaneously. BLAST [14] serves as a good example, which uses optimal instrumentation of loops and functions in order to achieve whole program control-flow attestation efficiently.
@@ 7.1.1 Hash Computation
Let us consider the above example once more. One solution to reduce the memory requirements on $\mathcal{P}$ is to make a single hash calculation over the collected control-flow transfers and return this (hashed) result to $\mathcal{V}$ after the CFG’s terminating node had been reached. This is method is, again, not widely due to the need to store every control-flow transfer before the hash function has been applied.
@@
Where $i$ denotes the node number, $H_{i}$ refers to the hash value generated at node $i$, and $h(\cdot)$ is a suitable hash function. $C_{i}$ is the control-flow event at $i$, which may be represented by a (src, dst) pair [12], an (entry, exit) pair [13], or a BBL ID [4]. $d$ represents the default value used when the first control-flow event when $i=0$, e.g. 00 in C-FLAT [4] or $h(0)$ in GACFA [9]. The main advantage of cumulative hash-based approach is that it does not require $\mathcal{P}$ to store all control-flow transfers. Once a round of calculation is completed, the corresponding control-flow transfers can be discarded. In Figure 16, we present a canonical example of a cumulative hash-based approach used by C-FLAT [4], demonstrating the hash values at different nodes of a target program’s CFG.
Figure 16: An example of cumulative hash calculation approach from C-FLAT [4]
@@
Standard cryptographic hash functions: BLAKE2 [85] is a common choice for CFA schemes targeted at embedded systems with limited computational resources [35, 13, 6, 43]. BLAKE2 is known to be significantly outperform ($\sim$50%) competing functions, such as Keccak and SHA-2, in raw execution time (cycles per byte) [86]. We note that SHA-256 [8, 9], SHA-512 [41] and SHA-3 [12, 34] have also been employed in existing work, albeit to a lesser extent. Worryingly, cryptographically broken algorithms have also been investigated by some schemes (MD5 [48, 41]).
@@
The space-efficient property of MSH functions can also pose some downsides. Namely, they do not consider the order of control-flow transfers in the set. It is possible that multiple legitimate control-flow paths have the same hash value, thus $\mathcal{V}$ loses contextual information control-flow path. Even if a hash value in the measurement result matches one in the database, $\mathcal{V}$ can determine only the number, not the order, of executed control-flow transfers. This could be fatal if an adversary finds a sequence of malicious control-flow transfers that produces the same MSH value as the legitimate sequence. We pose this issue as a research challenge.
@@ 7.1.2 Discussion
A segmented hashing approach has been used to reduce, but not eliminate, the effects of path explosion. Here, the target program is divided into multiple segments, and a separate hash computation is performed for each segment. Atrium [37] considers each loop in the target program as a separate, independent segment. Let us consider a toy target program containing five sequential loops, each of which may iterate up to 10 times. A separate hash computation is performed for each loop. $\mathcal{P}$ will generate six hash values to represent the entire control-flow path, one for each loop, plus one hash value to represent a control-flow path excluding all loops. Consequently, $\mathcal{V}$ must only store 51 legitimate hash values (10 for each of the five loops) and one legitimate for the control-flow path excluding loops, signfiicantly reducing the storage requirements versus a conventional hash-based approach (resulting in $\mathcal{V}$ storing 100,000 control-flow paths). Segmented hash computation has also been used to tackle conditional branches and recursive functions [13].
@@ 7.2 Other Path Log Representations
BLAST [14] uses Ball-Larus numbering to assign a numeric value to each edge of the target program’s CFG. Every possible path is marked by a unique integrity identifier,called pathID, which is the sum of the values of the edges included in the path. When a function call has finished execution, its pathID, target address of the function call, and return instruction are recorded and stored in shared memory between the REE and TEE. When a predetermined length has been reached, the log is sent to measurement module within the to perform a cumulative hash computation. BLAST can send two different measurement results. By default, $\mathcal{P}$ sends a single hash value as a measurement result to $\mathcal{V}$. If $\mathcal{V}$ cannot match the value in its database, it requests that $\mathcal{P}$ provides a sequence of tuples <func, pathID>. $\mathcal{P}$ then computes the Whole Program Path (WPP) representation [88] from the sequence and sends it to $\mathcal{V}$.
@@ 7.3 Whitelist-based Monitoring
One way to mitigate the problem caused long control-flow transfer sequences is using partial attestation reports. Here, $\mathcal{P}$ generates a mini-attestation report when specific requirements are meet—e.g. the configured length of the sequence is reached [34], and particular piece of code finishes [6]—in order to prevent memory exhaustion. Using partial attestation reports allows $\mathcal{V}$ to receive information about control-flow paths once certain portions of the target program has finished execution without waiting for it to terminate. Indeed, partial attestation reports can meet some scenarios with real-time requirements (see ARI [43]). Compression has also been used to minimise the measurement results in attestation reports; for example, ReCFA [58] used Zstandard [91], an off-the-shelf compression algorithm, to reduce the size of control-flow transfer sequences.
@@ 7.4 Abstract Execution
Abstract execution [92] is a method for monitoring authorised program states by simulating or mimicking the execution the proving program. This is principally performed on the side of $\mathcal{V}$ who, after receiving a set of program states from $\mathcal{P}$, checks whether it corresponds to a valid set of program states. In the context of CFA, this is similar to whitelist-based measurements. However, rather than representing control-flow transfers as $(src,dst)$, $\mathcal{V}$ simulates the execution of $T_{\mathcal{P}}$ by aggregating control-flow transfers using its CFG. The set of states from $\mathcal{P}$ is compared with legitimate, reconstructed paths in order to detect the presence of control-flow hijacking attacks. The representation of control-flow transfers more generally depends on different CFA proposals. One way is to represent indirect control-flow transfers by its $dst$ value and conditional branches using a binary encoding, where 0 = not taken and 1 = taken [34]. Another representation involves assigning dedicated IDs to target control-flow transfers [10].

### v3

@@ 5 Trust Anchors and Assumptions for Control-Flow Attestation
The trust anchor is commonly used to protect $\mathcal{P}$’s measurement procedure and any key material used in generating attestation reports. For some proposals, this requires that the trust anchor passes traditional static attestation and that it has dedicated memory to store material that only the anchor itself can access [41, 37, 13, 34]. For proposals that rely on TEEs, it is assumed that potentially applicable high-cost attacks cannot be performed by an adversary, such as hardware fault injections and side-channel attacks [51]. TEE-based proposals sometimes place trust in certain operating system components, such as kernel modules that initiate the communication with the TEE [38, 6]. In proposals where the kernel serves as $\mathcal{P}$’s trust anchor, it is somtimes assumed that a method for protecting cryptography keys is available, such as a TPM or a proprietary facility (see Intel MPK in [58]).
@@
  • Hardware-assisted systems use software whose security is underpinned by a hardware root of trust. TEE-based proposals are characterised as such, as well as proposals that rely on read-only memory (ROM), secure boot [45], and a TPM [13] in tandem with a normal operating system.
@@ 5.3 Central Processing Unit Features
Another common trust anchor is the use of custom or proprietary CPU features. These are predominately used to realise a measurement module that gathers control-flow information while the target program is being executed, i.e. at a machine code level. Such CFA schemes can be divided into two sub-categories. The first comprises schemes that employ custom extensions to CPU designs, e.g. open-source soft-core CPUs, to capture relevant control-flow events. The second contains proposals that rely on proprietary features of existing CPUs, such as tracing modules that track the history of instructions executed by the processor.
@@ 5.3.1 Custom Extensions
Figure 11: System architecture of LO-FAT [12].
@@
Atrium [37] (Figure 12) uses hardware extensions to capture instructions and memory addresses at the execution stage of the CPU’s pipeline. The extensions capture the control-flow changes taken by the target program on $\mathcal{P}$ while executing a challenge from $\mathcal{V}$. An instruction filter is used to consider only, and all, branch and jump instructions in addition to the current program counter values, and a separate loop encoder is used to track the execution of loops. A hashing engine is used to convert the executed instruction, loop and function IDs—corresponding to nodes of a CFG—into a sequence of cryptographic hashes produced in real-time during the program’s execution. These hashes are continuously stored in memory and used in the attestation report returned to $\mathcal{V}$. It is noted that Atrium has reduced on-chip memory requirements while also addressing TOCTOU attacks when compared with LO-FAT.
LiteHAX [34] is more complicated in nature, hooking three parts of the CPU’s pipeline: the decode, execute, and load/store phases (Figure 13). The proposal monitors all control-flow instructions and memory addresses of load/store instructions in order to address control- and data-flow attacks respectively. A similar hash-based measurement approach is used as Atrium [37] and LO-FAT [12]. However, the LiteHAX hash engine takes encoded control-flow instructions and load/store instructions, their memory addresses, and the current program counter value when calculating the hash values for the attestation report.
Figure 12: System architecture of Atrium [37].
@@ 5.3.2 Trace Modules
In recent years, CPU manufacturers have deployed capabilities that actively record control-flow information of programs under execution. This is known as program tracing whereby the CPU outputs encoded control-flow information into readable memory. Authorised applications can read this data, also known as a trace, to learn the control-flow path followed by a given program. The trace feature is supported by many contemporary CPU processors; for example, Intel CPUs support a tracing module from their 5th generation models, called Intel Processor Trace (PT) [64], which is designed for debugging pruposes. ARM CoreSight [65] is a comprehensive suite of debug and trace components that are embedded in ARM-based processors, providing developers with method to optimise and debug their applications on ARM architectures. The RISC-V specifications also describe a trace module; however, its availability varies by vendor due to RISC-V’s modular instruction set architecture [66, 67].
Figure 13: System architecture of LiteHAX [34].
@@ 5.4 Discrete Hardware Modules
Geden & Rasmussen [47] proposed a hardware security module (HSM) that collects control-flow information from the system bus (Figure 14). The HSM continuously takes control-flow events as input from the CPU, including the requested instruction address, opcode, and operand values over the system-on-chip’s control line and address and data buses. The collected events are compared with security policies defined in the HSM’s runtime integrity model for detecting code injection and re-use attacks. The use of continuous run-time data addresses TOCTOU attacks: a challenge-response protocol that measures events at only fixed points in time does not detect attackers that act transiently between those points. The proposal allows $\mathcal{V}$ to acquire a report from the HSM showing whether any security policy violations were detected; it also supports local verification whereby the module outputs a binary result to indicate if a target control-flow hijacking attack is detected. In scenarios where attacks are detected, the module outputs a binary result and provides auxiliary information, such as the last executed instruction.
Figure 14: The HSM used by Geden & Rasmussen [47].
@@ 6 Instrumentation Methods
Collecting control-flow data can be achieved through the process of instrumentation. This involves modifying the target program to add instructions that capture, and measure, control-flow events of interest. Dispatch instructions are used near target control-flow transfers or at the entry of BBLs to this end. When a target control-flow transfer is executed at run-time, the dispatch instructions call pass the transfer(s) to the measurement module for it to be measured. Execution is resumed when the measurement process has completed.
@@ 6.2 Binary Analysis and Rewriting
Recording every function call can impose significant performance overhead (see p. 3, [58]). To address this, ReCFA [58] derives a CFG from the target program using TypeArmor for performing call-site filtering, whereby CFG nodes are skipped where none of its predecessors has more than one successor. ReCFA uses Dyninst [81] for the static binary instrumentation of user-space programs on $\mathcal{P}$. Dyninst provides a set of libraries and utilities designed for the static (and dynamic) inspection and manipulation of executables. Dyninst enables users to insert, modify, or remove code in a program’s CFG without access to its source code. Intel’s MPK system is used to protect the integrity of instrumented data structures—the loop stack and path stack—used by control-flow events. The authors compare this to SCaRR [13] where control-flow edges are measured in kernel space, which has a “potential enormous context switching cost for the control-flow events on an order of magnitude in gigabytes” (p. 6, [58]). The authors combine kernel mode-based functions to preserve code integrity and MPK-protect data structures in order to uphold the integrity of the measurement procedure.
@@ 6.3 Dynamic Instrumentation
Dynamic instrumentation refers to the method of inserting, monitoring and verifying code into a program while it is running. This compares to instrumenting target programs using static analysis and manipulating the executable before execution (e.g. using binary rewriting). Dynamic methods rely on the insertion of probes or tracking mechanisms into a running program. These probes can monitor the control flow of the application in real-time, identifying any deviations from expected paths.
@@ 7 Measuring and Monitoring Control Flow
The measurement process is one of the key facets of a CFA scheme. It is at the core of $\mathcal{P}$ demonstrating to $\mathcal{V}$ that an authorised path/walk was followed through the target program’s CFG. For a given CFA scheme, the measurement method typically remains constant, regardless of the control-flow paths or the target program in question. However, these methods tend to vary significantly between CFA schemes, which we elaborate upon in this section.
@@
A CFA scheme may measure the whole program and partial elements, or multiple aspects of partial elements, simultaneously. BLAST [14] serves as a good example, which uses optimal instrumentation of loops and functions to achieve whole program control-flow attestation efficiently.
@@ 7.1.1 Hash Computation
Let us consider the above example once more. One solution to reduce the memory requirements on $\mathcal{P}$ is to make a single hash calculation over the collected control-flow transfers and return this (hashed) result to $\mathcal{V}$ after the CFG’s terminating node has been reached. This method is, again, not widely used due to the need to store every control-flow transfer before the hash function has been applied.
@@
Where $i$ denotes the node number, $H_{i}$ refers to the hash value generated at node $i$, and $h(\cdot)$ is a suitable hash function. $C_{i}$ is the control-flow event at $i$, which may be represented by a (src, dst) pair [12], an (entry, exit) pair [13], or a BBL ID [4]. $d$ represents the default value used with the first control-flow event when $i=0$, e.g. 00 in C-FLAT [4] or $h(0)$ in GACFA [9]. The main advantage of cumulative hash-based approach is that it does not require $\mathcal{P}$ to store all control-flow transfers. Once a round of calculation is completed, the corresponding control-flow transfers can be discarded. In Figure 16, we present a canonical example of a cumulative hash-based approach used by C-FLAT [4], demonstrating the hash values at different nodes of a target program’s CFG.
Figure 16: C-FLAT’s cumulative hash calculation approach [4].
@@
Standard cryptographic hash functions: BLAKE2 [85] is a common choice for CFA schemes targeted at embedded systems with limited computational resources [35, 13, 6, 43]. BLAKE2 is known to significantly outperform ($\sim$50%) competing functions, such as Keccak and SHA-2, in raw execution time (cycles per byte) [85]. We note that SHA-256 [8, 9], SHA-512 [41] and SHA-3 [12, 34] have also been employed in existing work, albeit to a lesser extent. Worryingly, cryptographically broken algorithms have also been investigated by some schemes (MD5 [49, 41]).
@@
The space-efficient property of MSH functions can also pose some downsides. Namely, they do not consider the order of control-flow transfers in the set. It is possible that multiple legitimate control-flow paths have the same hash value, thus $\mathcal{V}$ loses contextual control-flow path information. Even if a hash value in the measurement result matches one in the database, $\mathcal{V}$ can determine only the number, not the order, of executed control-flow transfers. This could be fatal if an adversary finds a sequence of malicious control-flow transfers that produces the same MSH value as the legitimate sequence. We pose this issue as a research challenge.
@@ 7.1.2 Discussion
A segmented hashing approach has been used to reduce, but not eliminate, the effects of path explosion. Here, the target program is divided into multiple segments, and a separate hash computation is performed for each segment. Atrium [37] considers each loop in the target program as a separate, independent segment. Let us consider a toy target program containing five sequential loops, each of which may iterate up to 10 times. A separate hash computation is performed for each loop. $\mathcal{P}$ will generate six hash values to represent the entire control-flow path, one for each loop, plus one hash value to represent a control-flow path excluding all loops. Consequently, $\mathcal{V}$ must only store 51 legitimate hash values (10 for each of the five loops) and one legitimate value for the control-flow path excluding loops, signfiicantly reducing the storage requirements versus a conventional hash-based approach (resulting in $\mathcal{V}$ storing 100,000 control-flow paths). Segmented hash computation has also been used to tackle conditional branches and recursive functions [13].
@@ 7.2 Other Path Log Representations
BLAST [14] uses Ball-Larus numbering to assign a numeric value to each edge of the target program’s CFG. Every possible path is marked by a unique integrity identifier—a path ID—which is the sum of the values of the edges included in the path. When a function call has finished execution, the path ID, target address of the function call, and return instruction are recorded and stored in shared memory between the REE and TEE. When a given length has been reached, the log is sent to measurement module within the TEE to compute a cumulative hash value. BLAST can send two different measurement results. By default, $\mathcal{P}$ sends a single hash value as a measurement result to $\mathcal{V}$. If $\mathcal{V}$ cannot match the value in its database, it requests that $\mathcal{P}$ provides a sequence of tuples <func, pathID>. $\mathcal{P}$ then computes the Whole Program Path (WPP) representation [87] from the sequence and sends it to $\mathcal{V}$.
@@ 7.3 Whitelist-based Monitoring
One way to mitigate the problem caused by long control-flow transfer sequences is using partial attestation reports. Here, $\mathcal{P}$ generates a mini-attestation report when specific requirements are met—e.g. the configured length of the sequence is reached [34], and particular piece of code finishes [6]—in order to prevent memory exhaustion. Using partial attestation reports allows $\mathcal{V}$ to receive information about control-flow paths once certain portions of the target program have finished execution without waiting for it to terminate. Indeed, partial attestation reports can meet some scenarios with real-time requirements (see ARI [43]). Compression has also been used to minimise the measurement results in attestation reports; for example, ReCFA [58] used Zstandard [90], an off-the-shelf compression algorithm, to reduce the size of control-flow transfer sequences.
@@ 7.4 Abstract Execution
Abstract execution [91] is a method for monitoring authorised program states by simulating or mimicking the execution of the proving program. This is principally performed on the side of $\mathcal{V}$ who, after receiving a set of program states from $\mathcal{P}$, checks whether it corresponds to a valid set of program states. In the context of CFA, this is similar to whitelist-based measurements. However, rather than representing control-flow transfers as $(src,dst)$, $\mathcal{V}$ simulates the execution of $T_{\mathcal{P}}$ by aggregating control-flow transfers using its CFG. The set of states from $\mathcal{P}$ is compared with legitimate, reconstructed paths in order to detect the presence of control-flow hijacking attacks. The representation of control-flow transfers more generally depends on different CFA proposals. One way is to represent indirect control-flow transfers by its $dst$ value and conditional branches using a binary encoding, where 0 = not taken and 1 = taken [34]. Another representation involves assigning dedicated IDs to target control-flow transfers [10].